Is the square root of 9 equal to -3i?
Answer by venugopalramana(3286)

NO NO >>>
-3i*-3i=+9*(i*i)=9*(-1)=-9
square root of a negative number only gives raise to complex numbers involving i..square root of a positrive number is always real and does not involve complex or imaginary number i
so square root of +9 could be +3 or -3